The lifetime of new engines T years has a continuous pdf f(t)={d/t2 if x>or equal1

{0,elsewhere

the value of the constant d,hence determine the mean and standard deviation T


1
Expert's answer
2021-07-16T11:56:17-0400


"\displaystyle\int_{-\infin}^{\infin}f(t)dt=\displaystyle\int_{1}^{\infin}\dfrac{d}{t^2}dt"

"=\lim\limits_{x\to \infin}[-\dfrac{d}{t}]\begin{matrix}
 x \\
 1
\end{matrix}=d=1"

"k=1"

"mean=\mu=E[T]=\displaystyle\int_{-\infin}^{\infin}tf(t)dt"

"=\displaystyle\int_{0}^{\infin}t(-\dfrac{1}{t^2})dt=\lim\limits_{x\to \infin}[-\ln t]\begin{matrix}
 x \\
 1
\end{matrix}=-\infin"


"E[T^2]=\displaystyle\int_{-\infin}^{\infin}t^2f(t)dt=\displaystyle\int_{0}^{\infin}t^2(-\dfrac{1}{t^2})dt"

"=\lim\limits_{x\to \infin}[- t]\begin{matrix}
 x \\
 1
\end{matrix}=-\infin"


"Var(T)=E[T^2]-(E[T])^2=-\infin"